OpenStudy (anonymous):

PLEASE I NEED HELP...I'LL AWARD MEDAL. Find The Length Of Arc AC. When The Central Angle Is 120 degrees and the radius is 10. (Use pi = 3.14) (The Correct Answer Should Be Rounded To 2 decimal places)

OpenStudy (phi):

Did you give it a try yet?

OpenStudy (anonymous):

Here's The Picture

OpenStudy (anonymous):

Yes I did give a try but i got the wrong answer @phi

OpenStudy (anonymous):

so can you help me out

OpenStudy (phi):

what did you do?

OpenStudy (anonymous):

i used the formula

OpenStudy (phi):

please write it out.

OpenStudy (anonymous):

and the formula is n degrees over 360 x 2 x pi x radius

OpenStudy (anonymous):

this formula is from my program

OpenStudy (phi):

yes. now fill in the numbers

OpenStudy (anonymous):

yeah I did that and keep getting 20.94395102 and when i round it to 2 decimal places i get 20.94 and i enter that....it was wrong

OpenStudy (phi):

just fill in the numbers. Don't simplify

OpenStudy (phi):

n /360 x 2 x pi x radius or 2*pi*r * n/360 pi is 3.14 what is r ? what is n?

OpenStudy (anonymous):

R is 10 and n is 120

OpenStudy (phi):

aren't we doing arc AC?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

@phi ???

OpenStudy (phi):

what is n? (for arc AC)

OpenStudy (phi):

do you see the little "box" that means "right angle"

OpenStudy (anonymous):

yeah

OpenStudy (phi):

how many degrees in a right angle ?

OpenStudy (anonymous):

90

OpenStudy (phi):

ok, you now have all the numbers. what do you get ?

OpenStudy (anonymous):

so n is 90

OpenStudy (phi):

yes. the "central angle" is 90º

OpenStudy (anonymous):

so i got 15.70796327

OpenStudy (phi):

they want you to use pi=3.14

OpenStudy (anonymous):

i did

OpenStudy (phi):

no, you used pi with more digits (that is how you got all those decimals)

OpenStudy (anonymous):

so the answer is 15.7

OpenStudy (anonymous):

right??

OpenStudy (phi):

if you use 3.14 you get 2*10*3.14 * ¼ = 62.8/4 = 15.7 I would type in 15.70

OpenStudy (anonymous):

i typed 15.7 and i got it right

OpenStudy (phi):

ok

OpenStudy (anonymous):

THANX!!!! @phi

OpenStudy (phi):

notice if you use pi= 3.14159 you get 15.70796327 and when you round to 2 places you get 15.71 and the program will mark it wrong (it really isn't, but it wants you to use 3.14)

OpenStudy (anonymous):

ok thanx